OpenStudy (lilkg77):

what is the point-slope equation of the line with slope -12 that goes through the point (5,3)? A. y-3=-12(x-5) B. y+3=-12(x+5) C. y+3=12(x+5) D. y-3=12(x-5)

OpenStudy (xmissalycatx):

I think it is D but that is my opinion. Let's ask @ganeshie8 and @dan815 for some help! :)

OpenStudy (lilkg77):

Are you sure its D?

OpenStudy (xmissalycatx):

I am not 100% confident in my answer which is why I was saying we should ask someone else for a second opinion.

OpenStudy (dan815):

a line that passes through point(a,b) and slope m is given by y-b=m(x-a)

OpenStudy (dan815):

here your slope is -12 (a,b)=(5,3) so y-3=-12(x-5)

OpenStudy (hihi67):

its either a or b because -12 is the slope and those are the only choices with -12

OpenStudy (lilkg77):

Thanks @dan815 for helping me understand the question